True or False

If the function "f" is odd, then "\displaystyle{\int_{-1}^{1}f(x)dx=0}"


1
Expert's answer
2021-12-14T05:56:17-0500

"\text{The solution is } \textit{True}
\\
\text{From } \int\limits_{ - 1}^1 {f\left( x \right)} dx = 0. \text{If we let } x = - t \implies dx = - dt
\\
\Rightarrow \int\limits_{ - 1}^1 {f\left( x \right)} dx = \int\limits_1^{ - 1} {f\left( { - t} \right)} dt = \int\limits_{ - 1}^1 {f\left( { - x} \right)} dx
\\
\text{If } f \text{ is , that is } f\left( { - x} \right) = - f\left( x \right)
\\
\Rightarrow \int\limits_{ - 1}^1 {f\left( x \right)} dx + \int\limits_{ - 1}^1 {f\left( { - x} \right)} dx = 0
\\
\text{Which is } \textbf{\textit{True}}"